OpenStudy (anonymous):

The formula for the perimeter P of a rectangle with length l and width w is P = 2l + 2w. Which of the following is a formula for the length of a rectangle in terms of the perimeter and width?

OpenStudy (anonymous):

A. l= p-w/2 B. l= p+w/2 C. l= p-2w/2 D. l= p+2w/2

OpenStudy (anonymous):

Just some algebraic manipulation, that's all. Basically, you can "tamper" with the equation as long as you tamper both sides of it equally. So, I can add 5 to both sides, and it will still be the same. How would you move the P to the other side? P = 2l + 2w

OpenStudy (anonymous):

You can also divide both sides by a number or variable, as long as you're not dividing a zero.